OpenStudy (anonymous):

6( X - 2.5 ) ≥8 - 6 ( 3.5 + X ) I Got X ≥ 6 as an answer

OpenStudy (anonymous):

i don't think that is right

OpenStudy (anonymous):

@satellite73 Explanation

OpenStudy (anonymous):

multiply out on both sides using the distributive law collect terms

OpenStudy (anonymous):

\[6x-15\geq 8-21-6x\] is a start

OpenStudy (anonymous):

@satellite73 so i re-did the math and I got x ≥ 5.5

OpenStudy (anonymous):

hmm lets check

OpenStudy (anonymous):

\[6x-15\geq 8-21-6x\] \[6x-15\geq -13-6x\] first then add \(6x\) to both sides, giving \[12x-15\geq -13\] add 15 and get \[12x\geq 2\]

OpenStudy (anonymous):

so \[x\geq \frac{2}{12}=\frac{1}{6}\]